UID Iz 2 47? 33. Rewrite the equation x2 - 6x - 11 = 0 in graphing form by completing the square. Give the location of the ve questions 33 and 34
Answer #1

Solution! 33 Given x²-6x-11=0 2 Converting the ginen equation in graphing from x² - 6x-11-0 = (22 – 2x3x + 33-32-11 - 0 (x-3).. girm by where a = initial value Exponentia fonesia -x 2. a. 62 and 64 .-3 E a b 42=ax Solution : 34 Exponential function G